OpenStudy (anonymous):

5^x+1 = 24 exponential equations

OpenStudy (anonymous):

same as before \[x+1=\frac{\ln(14)}{\ln(5)}\]

OpenStudy (anonymous):

i got the answer 1

OpenStudy (anonymous):

is that right?